OpenStudy (anonymous):

The distance from one corner to the opposite corner of a square playground is 96 ft. To the nearest foot, how long is each side of the playground?

OpenStudy (anonymous):

Diagonal of a square is \(\sqrt{2} \times \) side.

OpenStudy (anonymous):



OpenStudy (anonymous):

So it should be \( 48 \sqrt{2} \) ft

OpenStudy (anonymous):

\[= 67.8823 \approx 68 ft \]

OpenStudy (anonymous):

Then x equals roughly 68ft per side, huh?

OpenStudy (anonymous):

Yes that's right.

OpenStudy (anonymous):

so to the nearest foot would be 68ft

OpenStudy (anonymous):

Thank you everybody!

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

Fool is awesome